OpenStudy (anonymous):

what does it mean l 4-3xl-4<11

OpenStudy (sgadi):

|4-3x|<15 -(4-3x)<15 or (4-3x)<15 (4-3x)>-15 or (4-3x)<15 -3x>-19 or -3x<11 3x<19 or 3x>-11 x<19/3 or x>-11/3 or we can say -11/3<x<19/3

OpenStudy (anonymous):

\[\left| 4-3x \right|<11+4=\left| 4-3x \right|<15=4-3x=15 or 4-3x=-15 solving 4-3x=15 we have x=-11/3,solving 4-3x=-15 we have x=19/3\]

OpenStudy (anonymous):

some problem occurred in equation editor,solution is the same as sgadi did.